OpenStudy (nathanjhw):

If you added 100g of hot water (43 degrees Celsius) to 100g of room temperature water (21 degrees Celsius) approximately how long would it heat up for and how hot would it get? P.S. I just need generalizations.

OpenStudy (aaronq):

"how long would it heat up for?" what do you mean by this? the general equation for these processes is: \(q=m*C_p*\Delta T\) since they have the same specific heat capacity (both substances are water), the equation is: \(q_{hot}=-q_{cold}\rightarrow m_{hot}*\Delta T=-m_{cold}*\Delta T\) \(m_{hot}*(T_f-T_i)=-m_{cold}*(T_f-T_i)\) \(m_{hot}*(T_f-43^oC)=-m_{cold}*(T_f-21^oC)\) from there you can find the final temperature of the mixture

OpenStudy (nathanjhw):

I just want to know approximately how long something heats up for before getting cold.

OpenStudy (nathanjhw):

@aaronq

OpenStudy (nathanjhw):

For example, if I added hot coffee to room temperature cream how long would that cream heat up for before hitting a plateau and cooling down?

OpenStudy (aaronq):

ohh, use newtons law of cooling. you need the ambient temperature though